OpenStudy (anonymous):

What is the value of the expression 3–4?

OpenStudy (anonymous):

You mean 3^-4?

OpenStudy (anonymous):

Do you know negative numbers

jimthompson5910 (jim_thompson5910):

Think of it like this you have $3 but you want to buy something that is $4 will you be able to buy it? Yes or no? If not, then how much do you have to borrow to buy it?

OpenStudy (anonymous):

Yeah please help me i'm in 7th grade and i need help

OpenStudy (anonymous):

1 dollar

OpenStudy (anonymous):

Thats not an option

OpenStudy (anonymous):

Sorry I meant -1

OpenStudy (anonymous):

\[\Huge 3-4 = -1 \]

OpenStudy (anonymous):

Your Textbook is obviously incorrect if this is you correct question I am sure @jim_thompson5910 agrees with me

OpenStudy (anonymous):

its 3 to the -4th power

OpenStudy (anonymous):

?

OpenStudy (anonymous):

So??

OpenStudy (anonymous):

I a m c o n f u s e d

OpenStudy (anonymous):

3 to the negative 4th power isn't even a whole number. What are your answer choices?

OpenStudy (anonymous):

@Drakula what are your answer choices?

OpenStudy (anonymous):

I just asked it it was 3^-4! o_o

OpenStudy (anonymous):

3^-4 = 1/3^4 = 1/81